OpenStudy (anonymous):

Need help!! working on this for a while!! Aluminum burns in liquid bromine according to the following equation: 2 Al(s) + 3 Br2(l) = 2 AlBr3(s) After reacting 20.0 g of aluminum, you isolate 87.4 g of aluminum bromide (AlBr3). a). What was the theoretical yield for the reaction? b). What was the percent yield for the reaction?

OpenStudy (zpupster):

1) Balance the equation: 2 Al(s) + 3 Br2(l) = 2 AlBr3(s) 2) Find the limiting reagent : divide by Molecular wt. Al=26.98g/mol 20g/(26.98 g/mol) = .74 moles 3)Find the theoretical yield :multiply the amount of moles of the limiting reagent by the ratio of the limiting reagent and the synthesized product and by the molecular weight of the product. 2 moles Al -->2 moles Al Br3 .74 moles of Al (2molesAlBr3/2moles Al)(266.8g/mol AlBr3) = 197.34 grams 4) Actual yield : given was 87.4g 5)Percentage Yield = (Mass of Actual Yield / Mass of Theoretical Yield) ( x 100%) 87.4/197.34 =44.29%